# Narrative Scaffolding: Transforming Data-Driven Sensemaking Through Narrative-First Exploration

### v1

@@ 3.1.2. Follow-up Interview ($N=5$) and Challenge Formation
The interviews revealed rich insights which directly informed our design goals. Open-ended survey responses and interview transcripts were analyzed using inductive thematic analysis (Nowell et al., 2017). Table 1 summarizes the four identified challenges (C1–C4), observed behaviors (B1–B9), and corresponding design goals (DG1–DG4).
@@ 3.2.1. Challenge 1: Unsupported Transition from Intent to Evidence
Prior work shows that analysts often begin exploration with tentative or underspecified intentions, yet most analytic systems demand premature precision or rigid query formulation (Pirolli and Card, 2005; Heer et al., 2008; Ragan et al., 2015). In our survey, 39 of 48 participants reported that their early exploration goals lacked detail (B1). $F_{12}$ reflected that ”It gives me something too specific when I’m still just exploring”. This mismatch limits how users pursue their curiosity or flexibly define intentions.
@@
To manage the situation, participants (36 of 48) offloaded their reasoning into external note-taking tools, such as slides or documents (B3). $F_{6}$ noted, ”I’ll write down an idea like ‘maybe this drop is due to [factor],’ but that note just lives separately from the data.” These workarounds highlight how reasoning fragments drift apart from evidence, creating temporal degradation. As $F^{\prime}_{2}$ reflected, ”I usually remember my conclusions, but not the reasoning that led me there.” This illustrates how outcomes are retained while the reasoning behind them fades, making it harder to revisit or refine past analyses.
@@ 3.2.2. Challenge 2: Unmanaged Inquiry and Reflective Continuity
Even when participants attempted to revisit their earlier reasoning, the process was limited and inconsistent. 32 of 48 participants reported rarely reviewing past conclusions (B5). $F’_{5}$ described, ”I sometimes go back to see if my conclusion still makes sense, but it’s hard to remember the context of why I thought it.” Without preserved context, retrospective reflection becomes fragmented and difficult to sustain (Hullman, 2022; Kale et al., 2022).
Finally, reasoning paths often ended unexpectedly as participants shifted focus. 25 of 48 participants acknowledged abandoning questions without resolution (B6). As $F’_{2}$ noted, ”I left one question midway, but I never went back to see if it was still valid.” This lack of continuity results in abandoned questions and forgotten assumptions, making it difficult to trace how reasoning evolved.
@@ 3.2.3. Challenge 3: Disrupted Divergence in Exploratory Reasoning
Most analytic tools record actions in a single linear sequence (Gotz and Zhou, 2009; Gratzl et al., 2016), but reasoning often branches as analysts test alternatives (Pirolli and Card, 2005; North et al., 2011). Participants (29 of 48) reported restarting or duplicating their workspace when their thinking shifted (B7). $F_{21}$ explained, ”If I want to try a new idea, I just duplicate the current exploration path and restart another, changing the original feels risky.” Restarting in this way creates a break in continuity: earlier interpretations are severed from the new line of thought, forcing participants to ”manage multiple versions manually” ($F_{43}$).
Prior work has shown that a narrative-driven approach can inadvertently reinforce existing assumptions when alternative explanations are not explored (Nickerson, 1998; Li et al., 2025; Wall et al., 2017). Participants (33 of 48) emphasized the importance of considering alternative explanations to maintain a balanced approach to their reasoning. However, a further 23 of 48 described difficulty keeping track of multiple possible directions when their reasoning diverged (B8). $F’_{4}$ reflected, ”When I try two different parallel explanations, I usually lose track of one of them.” Without support to manage alternatives side by side, users struggled to weigh competing possibilities, which could limit consideration of disconfirming evidence.
@@ 4. Narrative Scaffolding Framework
4.1. Framework Stages and Theoretical Grounding
To connect our established design goals (DG1-DG4) to actual analytic practice — and use it to ground the design of our system (presented in Sec 5), we propose a conceptual model that captures how narrative construction unfolds through distinct stages of interpretive reasoning. This model provides a structured view of the cognitive processes our system is designed to support. The model was adapted from two established perspectives on analytic reasoning: the Sensemaking Loop (Pirolli and Card, 2005), which describes an iterative cycle between searching for information and organizing it into mental structures, and the Data–Frame model (Klein et al., 2007a), which emphasizes how analysts construct, test, and revise interpretive frames in response to new evidence. We also draw on external cognition theory (Scaife and Rogers, 1996) and narrative psychology (Murray, 2003) to ground how each stage supports reasoning. Building on these foundations, as shown in Fig. 2, our adaptation repositions narrative construction as the primary reasoning interface, and maps four stages to our design goals.
@@ 4.1. Framework Stages and Theoretical Grounding
Each stage operationalizes specific theoretical constructs. Narrative Development draws on external cognition theory: by externalizing assumptions as written text, analysts transform fleeting thoughts into manipulable artifacts that can be inspected and revised (Scaife and Rogers, 1996; Liu and Stasko, 2010). Investigation operationalizes the Data-Frame model’s emphasis on testing interpretive frames against evidence, where claims and data mutually reshape one another (Klein et al., 2007a). Reflection embodies the Sensemaking Loop’s iterative cycling between bottom-up evidence gathering and top-down schema revision (Pirolli and Card, 2005). Integration draws on narrative psychology’s insight that constructing coherent narratives is itself a mode of reasoning, not merely a post-hoc communication step (Murray, 2003; Bruner, 1991).
@@ 5.3.1. Sentence-to-View Linking
Figure 5. Sentence-level reflection allows analysts to revisit specific narrative statements in context, linking individual sentences to visual evidence for immediate, localized reasoning.
@@ 5.3.2. Insight Provenance Timeline
The insight timeline captures how insights emerge and evolve as users write, recording each update as a node (Figure 6). We adapted the Data–Frame taxonomy (Klein et al., 2007a; Yi et al., 2008) to classify insights into four relationship types describing how claims shift the trajectory of reasoning: provide overview, adjust, detect pattern, and match mental model. The timeline is organized as a tree structure, preserving both the current path and alternative branches. At this level, reflection focuses on how a user’s thinking shifts over time, helping users answer questions such as ”How did I arrive at this claim?” or ”Has my hypothesis drifted?”. By visualizing the trajectory of reasoning as a tree, this design provides a navigable record that allows users to revisit and recover the precise state of thought at any point in their exploration.
@@ 5.3.3. Inquiry Board
The inquiry board captures how inquiries unfold over time (Figure 7). Inquiries are derived directly from the narrative: when users write hypotheses or intentions to explore, these statements are formalized as issues, while claims provide potential answers. Because exploration is ongoing, many inquiries may remain unanswered or give rise to new ones as paths branch. The board categorizes all current inquiries into three statuses: open, resolved, or stalled, and arranges them in an IBIS graph (Conklin and Begeman, 1988; Kunz and Rittel, 1970) to show how issues relate to one another. At this level, reflection shifts from local claims to the broader organization of reasoning. The inquiry board enables users to revisit unresolved issues, re-engage with abandoned threads, and assess whether current claims have sufficient support, preventing forgotten questions from fading silently out of the reasoning process.
Figure 7. Inquiry-level reflection synthesizes across insights, reviewing how questions evolve and progress over time.
@@ 5.6. Implementation Details
Narrative Scaffolding is implemented as a web application with a Python Flask backend and React Next.js frontend. We employ a chain-of-thought (Wei et al., 2022) prompting strategy to make LLM reasoning more explicit and context-aware. GPT-5 was used to handle higher-level narrative generation tasks that require longer context windows and complex reasoning for storyline development, ensuring faithful mappings between user intent and chart configurations. GPT-4o-mini was used to manage lightweight operations where responsiveness is critical, such as insight extraction, inquiry summarization, and reflective nudges. This approach was chosen based on empirical testing that balanced consistency, speed, and task-specific performance.
Data visualizations utilize Vega-Lite for chart rendering and Leaflet.js for interactive maps. Text-to-SQL functionality leverages VannaAI (Vanna.AI Team, 2024) with OpenAI embedding and few-shot training for query consistency. The system is designed to be model-agnostic, allowing for easy substitution of LLMs as capabilities evolve. Complete prompt templates and configuration details are provided in supplementary materials.
@@ 4. RQ4
    [Story Explainability] How does NS support structured and defensible narratives (DG4), making users’ data stories more explainable to others?
@@ 6.1.1. Baseline and Apparatus
In addition to our NS system, we implemented a Baseline for comparison. The Baseline is a web-based analytics workbench that replicates typical Power BI workflows (Microsoft, 2024b), supporting natural-language querying and visualization generation within a traditional dashboard interface. Participants could issue queries, apply filters, generate charts, and export findings using conventional point-and-click interactions supplemented by conversational AI assistance. To ensure a fair comparison focused on workflow design rather than technical capabilities, both systems operated on the same data schema, text-to-SQL translation, and semantic view generation pipeline (5.2.1). For the final integration stage, we exported tabular data from generated visuals via the Power BI REST API to generate narrative summaries of trends, comparisons, and anomalies (Microsoft, 2024a).
@@ 6.1.2. Method
[Procedure]: We adopted a within-subject design where each participant used both NS and Baseline, with tool order counterbalanced to mitigate learning effects (Figure  8). Sessions began with informed consent, demographics collection, and system introduction. For each system, participants received a tutorial, completed the assigned task while thinking aloud (Eccles and Arsal, 2017) (explicitly verbalizing reflection, exploration, and insights), and completed a post-task survey evaluating usability, workload, and experience. After both systems, participants completed a final comparison survey reflecting on workflow differences. Each session lasted approximately 90 minutes, and participants received \$20 USD compensation upon completion. Study instructions and survey instruments are provided in the supplementary materials.
@@ 6.1.3. Measures and Analysis
We collected both perception and interaction data to evaluate the two systems comprehensively. We analyzed post-task survey responses that included 7-point Likert scale questions on workflow efficiency, exploration, reflection, and explainability. To assess usability and workload, we incorporated the SUS (Brooke et al., 1996) and NASA-TLX (Hart, 2006), alongside self-assessments of the clarity and defensibility of participants’ final narratives. Interaction logs captured differences in analysis patterns, including query diversity, revisions, and visualizations created, with temporal measures normalized across sessions. In addition, we recorded think-aloud protocols during task sessions to capture participants’ reasoning processes in situ. These transcripts were then thematically analyzed to account for the reflection and exploration process. We compared survey responses between the Baseline and NS conditions using the Wilcoxon signed-rank test, given the ordinal nature of Likert-scale data. For continuous temporal measures, we applied Welch’s t-test when the normality assumptions were met and the Mann–Whitney U test otherwise. To control for false discovery due to multiple comparisons, we applied the Benjamini–Hochberg procedure (Benjamini and Hochberg, 1995). We report Cohen’s d effect sizes for all parametric comparisons to indicate the magnitude of practical differences between conditions.
@@ 7. Results
We present an analysis of survey responses and participant interactions ($P_{1}-P_{20}$) with both conditions. Our results show that NS enabled participants to generate insights more efficiently, engage in deeper and more continuous reflection, explore broader analytical alternatives, and produce more defensible narratives compared to the baseline condition. NS received an average System Usability Score (SUS) of 79.8 (Figure 9), demonstrated high usability (Excellent according to (Bangor et al., 2009) and passing the cutoff score, 68, for production by a large margin). $P_{11}$ described the system as feeling ”like a reliable companion” during analysis.
@@ 7.1.3. Participants Found Direction and Generated Structured Output More Easily
NS also significantly improved participants’ ability to turn exploration into structured output (Figure 11, B) (NS: M = 5.45, SD = 0.89; Baseline: M = 3.50, SD = 1.06, p = .002). $P_{7}$ found that ”turning exploration into a report feels almost effortless,” and $P_{17}$ highlighted how the system ”transforms explored branches into a structured story with one click.” These ratings align with the behavioural efficiency measures, indicating that participants both experienced and recognized the workflow improvements provided by NS, supporting DG1’s goal of fluid transitions from vague inquiries to evidence-grounded insights.
@@ 7.3.1. Participants Explored More Alternatives and Factors
With NS, participants considered significantly more alternatives during their analysis (M = 8.75, SD = 3.03, p ¡ 0.001) compared to Baseline participants (M = 3.25, SD = 1.58). They also explored a greater breadth of available factors, examining on average 7.0 out of 8 possible factors (SD = 0.9), compared to Baseline participants who explored 5.3 factors (SD = 1.4). As $P_{8}$ explained, “With NS I felt like I could easily try different takes on the data, instead of sticking to one angle.”
@@
$P_{15}$ emphasized how NS encouraged breadth across factors: “The system had nudged me to look at other dimensions when I reflected, like community well-being alongside cost.” Together, these findings suggest that NS not only increased the number of alternatives participants explored but also helped them consider broader coverage of available factors, creating more opportunities to surface contrasting perspectives.
@@ 7.3.2. Participants Integrated More Evidence into Decision-Making
Beyond exploring more alternatives, NS participants integrated more of their exploration into final decisions. They incorporated significantly more factors into their decision-making process (M = 3.8, SD = 0.7) compared to Baseline participants (M = 2.5, SD = 0.7). This suggests that NS participants not only explored more broadly but also utilized their exploration more meaningfully. Participants emphasized that the more exhaustive exploration in NS translated into having more actionable options to weigh. As $P_{11}$ explained, “Because I had looked at more angles, I actually had choices, I could go in a few different directions instead of just one.” Similarly, $P_{6}$ elaborated, “Having more alternatives in [NS] meant I didn’t have to stick with my first idea […] I could line up different options and actually remember how they compared, which made it easier to choose from.”
@@
7.4. [RQ4]: Story Explainability
Participants using NS reported significantly greater confidence in defending the reasoning behind their insights (M = 5.90, SD = 1.25) compared to Baseline (M = 3.15, SD = 1.60, p ¡ .001). Participants described how timelines and provenance features supported justification. $P_{2}$ explained, “The timelines show the whole process which guarantees transparency and provides relevant evidence.” $P_{13}$ noted, “Since the insights were linked to evidence and captured in the timeline, it is pretty straightforward to justify my reasoning.” Similarly, $P_{17}$ emphasized, “Since every captured insight is automatically linked back to its supporting data and visualizations, I would always have clear evidence to justify my claims.” $P_{20}$ summarized this benefit as, “I felt more confident that I could defend my choices because I had pulled in evidence from different places, not just one metric.”
When asked about presenting conclusions to others, participants also reported higher average confidence with NS (M = 5.95, SD = 1.00) than with Baseline (M = 5.15, SD = 1.60). However, this difference was not statistically significant (p = .112). Participants ($N=13$) rated themselves as confident presenting with NS, citing features such as organized evidence, visible branches, and insight timelines, which serve as valuable sources of support. For instance, $P_{1}$ stated, “With all evidence organized, branches visible, and a polished story generated, I’d feel confident explaining my analysis to anyone.” Similarly, $P_{10}$ noted, “Every insight is backed by reasoning, and the story mode is also helpful; therefore, it will aid in justifying my recommendations.”
Among these same participants, ($N=9$) also rated themselves as confident presenting with Baseline. In these cases, confidence stemmed from their own skills rather than the tool itself: $P_{1}$ remarked, “I feel really confident because once I make a presentation, I will be able to address the weak points in my insights and conclusions.” Similarly, $P_{12}$ explained, “I’d be comfortable with the recommendations themselves, but less confident explaining the full path I took to get there since some of that context lives outside the tool.”
@@
Experts explained that lower ratings often reflected a failure to capture implied intent. While generated dashboards were usually relevant to the surface wording of the prompt, they sometimes missed the deeper analytical angle that vague phrasing implied. For example, $E_{3}$ noted a prompt about “some neighbourhoods are unexpectedly safe” suggested an outlier-focused analysis, but the system produced generic averages and correlations. Similarly, $E_{5}$ explained that when asked about ”travel growth differences between countries” the system produced growth rate visualizations that ”It showed similar percentages across countries, but missed that 10% growth could mean millions in one place and only thousands in another.” $E_{2}$ described such cases as “technically correct but superficial”. This implies that while the majority of outputs were judged as good matches to the prompts, a smaller subset still fell short by failing to respond to the implied meaning behind vague inquiries.

### v2

@@ 3.1.2. Follow-up Interview ($N=5$) and Challenge Formation
The interviews revealed rich insights which directly informed our design goals. Open-ended survey responses and interview transcripts were analyzed using inductive thematic analysis (nowell2017thematic). Table 1 summarizes the four identified challenges (C1–C4), observed behaviors (B1–B9), and corresponding design goals (DG1–DG4).
@@ 3.2.1. Challenge 1: Unsupported Transition from Intent to Evidence
Prior work shows that analysts often begin exploration with tentative or underspecified intentions, yet most analytic systems demand premature precision or rigid query formulation (pirolli2005sensemaking; heer2008graphical; ragan2015characterizing). In our survey, 39 of 48 participants reported that their early exploration goals lacked detail (B1). $F_{12}$ reflected that ”It gives me something too specific when I’m still just exploring”. This mismatch limits how users pursue their curiosity or flexibly define intentions.
@@
To manage the situation, participants (36 of 48) offloaded their reasoning into external note-taking tools, such as slides or documents (B3). $F_{6}$ noted, ”I’ll write down an idea like ’maybe this drop is due to [factor],’ but that note just lives separately from the data.” These workarounds highlight how reasoning fragments drift apart from evidence, creating temporal degradation. As $F^{\prime}_{2}$ reflected, ”I usually remember my conclusions, but not the reasoning that led me there.” This illustrates how outcomes are retained while the reasoning behind them fades, making it harder to revisit or refine past analyses.
@@ 3.2.2. Challenge 2: Unmanaged Inquiry and Reflective Continuity
Even when participants attempted to revisit their earlier reasoning, the process was limited and inconsistent. 32 of 48 participants reported rarely reviewing past conclusions (B5). $F^{\prime}_{5}$ described, ”I sometimes go back to see if my conclusion still makes sense, but it’s hard to remember the context of why I thought it.” Without preserved context, retrospective reflection becomes fragmented and difficult to sustain (hullman2022judgment; kale2022discern).
Finally, reasoning paths often ended unexpectedly as participants shifted focus. 25 of 48 participants acknowledged abandoning questions without resolution (B6). As $F^{\prime}_{2}$ noted, ”I left one question midway, but I never went back to see if it was still valid.” This lack of continuity results in abandoned questions and forgotten assumptions, making it difficult to trace how reasoning evolved.
@@ 3.2.3. Challenge 3: Disrupted Divergence in Exploratory Reasoning
Most analytic tools record actions in a single linear sequence (gotz2009characterizing; gratzl2016visual), but reasoning often branches as analysts test alternatives (pirolli2005sensemaking; north2011analytic). Participants (29 of 48) reported restarting or duplicating their workspace when their thinking shifted (B7). $F_{21}$ explained, ”If I want to try a new idea, I just duplicate the current exploration path and restart another, changing the original feels risky.” Restarting in this way creates a break in continuity: earlier interpretations are severed from the new line of thought, forcing participants to ”manage multiple versions manually” ($F_{43}$).
Prior work has shown that a narrative-driven approach can inadvertently reinforce existing assumptions when alternative explanations are not explored (nickerson1998confirmation; confirmation2025; wall2017warning). Participants (33 of 48) emphasized the importance of considering alternative explanations to maintain a balanced approach to their reasoning. However, a further 23 of 48 described difficulty keeping track of multiple possible directions when their reasoning diverged (B8). $F^{\prime}_{4}$ reflected, ”When I try two different parallel explanations, I usually lose track of one of them.” Without support to manage alternatives side by side, users struggled to weigh competing possibilities, which could limit consideration of disconfirming evidence.
@@ 4. Narrative Scaffolding Framework
4.1. Stages and Theoretical Grounding
To connect our established design goals (DG1-DG4) to actual analytic practice — and use it to ground the design of our system (presented in Sec 5), we propose a conceptual model that captures how narrative construction unfolds through distinct stages of interpretive reasoning. This model provides a structured view of the cognitive processes our system is designed to support. The model was adapted from two established perspectives on analytic reasoning: the Sensemaking Loop (pirolli2005sensemaking), which describes an iterative cycle between searching for information and organizing it into mental structures, and the Data–Frame model (klein2007data), which emphasizes how analysts construct, test, and revise interpretive frames in response to new evidence. We also draw on external cognition theory (scaife1996external) and narrative psychology (murray2003narrative) to ground how each stage supports reasoning. Building on these foundations, as shown in Fig. 2, our adaptation repositions narrative construction as the primary reasoning interface, and maps four stages to our design goals.
@@ 4.1. Stages and Theoretical Grounding
Each stage operationalizes specific theoretical constructs. Narrative Development draws on external cognition theory: by externalizing assumptions as written text, analysts transform fleeting thoughts into manipulable artifacts that can be inspected and revised (scaife1996external; liu2010mental). Investigation operationalizes the Data-Frame model’s emphasis on testing interpretive frames against evidence, where claims and data mutually reshape one another (klein2007data). Reflection embodies the Sensemaking Loop’s iterative cycling between bottom-up evidence gathering and top-down schema revision (pirolli2005sensemaking). Integration draws on narrative psychology’s insight that constructing coherent narratives is itself a mode of reasoning, not merely a post-hoc communication step (murray2003narrative; bruner1991narrative).
@@ 5.3.1. Sentence-to-View Linking
Figure 5. Sentence-level reflection allows analysts to revisit specific narrative statements in context, linking individual sentences to visual evidence for localized reasoning.
@@ 5.3.2. Insight Provenance Timeline
The insight timeline captures how insights emerge and evolve as users write, recording each update as a node (Figure 7). We adapted the Data–Frame taxonomy (klein2007data; yi_insights_2008) to classify insights into four relationship types describing how claims shift the trajectory of reasoning: provide overview, adjust, detect pattern, and match mental model. The timeline is organized as a tree structure, preserving both the current path and alternative branches. At this level, reflection focuses on how a user’s thinking shifts over time, helping users answer questions such as ”How did I arrive at this claim?” or ”Has my hypothesis drifted?”. By visualizing the trajectory of reasoning as a tree, this design provides a navigable record that allows users to revisit and recover the precise state of thought at any point in their exploration.
@@
Figure 7. Inquiry-level reflection synthesizes across insights, reviewing and linking how questions evolve and progress over time.
@@ 5.3.3. Inquiry Board
The inquiry board captures how inquiries unfold over time (Figure 7). Inquiries are derived directly from the narrative: when users write hypotheses or intentions to explore, these statements are formalized as issues, while claims provide potential answers. Because exploration is ongoing, many inquiries may remain unanswered or give rise to new ones as paths branch. The board categorizes all current inquiries into three statuses: open, resolved, or stalled, and arranges them in an IBIS graph (conklin1988gibis; kunz1970issues) to show how issues relate to one another. At this level, reflection shifts from local claims to the broader organization of reasoning. The inquiry board enables users to revisit unresolved issues, re-engage with abandoned threads, and assess whether current claims have sufficient support, preventing forgotten questions from fading silently out of the reasoning process.
@@ 5.6. Implementation Details
Narrative Scaffolding is implemented as a web application with a Python Flask backend and React Next.js frontend. We employ a chain-of-thought (wei2022chain) prompting strategy to make LLM reasoning more explicit and context-aware. GPT-5 was used to handle higher-level narrative generation tasks that require longer context windows and complex reasoning for storyline development, ensuring faithful mappings between user intent and chart configurations. GPT-4o-mini was used to manage lightweight operations where responsiveness is critical, such as insight extraction, inquiry summarization, and reflective nudges. This approach was chosen based on empirical testing that balanced consistency, speed, and task-specific performance.
Data visualizations utilize Vega-Lite for chart rendering and Leaflet.js for interactive maps. Text-to-SQL functionality leverages VannaAI (vannaAI) with OpenAI embedding and few-shot training for query consistency. The system is designed to be model-agnostic, allowing for easy substitution of LLMs as capabilities evolve. Complete prompt templates and configuration details are provided in supplementary materials.
@@ 4. RQ4
    [Reasoning Traceability] How does NS support traceable reasoning (DG4), making the analytical process behind insights more visible and defensible?
@@ 6.1.1. Baseline and Apparatus
In addition to our NS system, we implemented a Baseline for comparison. The Baseline is a web-based analytics workbench that replicates typical Power BI workflows (microsoftCopilotIntro), supporting natural-language querying and visualization generation within a traditional dashboard interface. Participants could issue queries, apply filters, generate charts, and export findings using conventional point-and-click interactions supplemented by conversational AI assistance. To ensure a fair comparison focused on workflow design rather than technical capabilities, both systems operated on the same data schema, text-to-SQL translation, and semantic view generation pipeline (5.2.1). For the final integration stage, we exported tabular data from generated visuals via the Power BI REST API to generate narrative summaries of trends, comparisons, and anomalies (microsoftPowerBINarrative).
@@ 6.1.2. Method
[Procedure]: We adopted a within-subject design where each participant used both NS and Baseline, with tool order counterbalanced to mitigate learning effects (Figure  9). Sessions began with informed consent, demographics collection, and system introduction. For each system, participants received a tutorial, completed the assigned task while thinking aloud (eccles2017think) (explicitly verbalizing reflection, exploration, and insights), and completed a post-task survey evaluating usability, workload, and experience. After both systems, participants completed a final comparison survey reflecting on workflow differences. Each session lasted approximately 90 minutes, and participants received \$20 USD compensation upon completion. Study instructions and survey instruments are provided in the supplementary materials.
@@ 6.1.3. Measures and Analysis
We collected both perception and interaction data to evaluate the two systems comprehensively. We analyzed post-task survey responses that included 7-point Likert scale questions on workflow efficiency, exploration, reflection, and explainability. To assess usability and workload, we incorporated the SUS (brooke1996sus) and NASA-TLX (hart2006nasa), alongside self-assessments of the clarity and defensibility of participants’ final narratives. Interaction logs captured differences in analysis patterns, including query diversity, revisions, and visualizations created, with temporal measures normalized across sessions. In addition, we recorded think-aloud protocols during task sessions to capture participants’ reasoning processes in situ. These transcripts were then thematically analyzed to account for the reflection and exploration process. We compared survey responses between the Baseline and NS conditions using the Wilcoxon signed-rank test, given the ordinal nature of Likert-scale data. For continuous temporal measures, we applied Welch’s t-test when the normality assumptions were met and the Mann–Whitney U test otherwise. To control for false discovery due to multiple comparisons, we applied the Benjamini–Hochberg procedure (benjamini1995controlling). We report Cohen’s d effect sizes for all parametric comparisons to indicate the magnitude of practical differences between conditions.
@@ 7. Results
We present an analysis of survey responses and participant interactions ($P_{1}-P_{20}$) with both conditions. Our results show that NS enabled participants to generate insights more efficiently, engage in deeper and more continuous reflection, explore broader analytical alternatives, and produce more defensible narratives compared to the baseline condition. NS received an average System Usability Score (SUS) of 79.8 (Figure 9), demonstrated high usability (Excellent according to (bangor2009determining) and passing the cutoff score, 68, for production by a large margin). $P_{11}$ described the system as feeling ”like a reliable companion” during analysis.
@@ 7.1.3. Participants Found Direction and Generated Structured Output More Easily
NS also significantly improved participants’ ability to turn exploration into structured output (Figure 11, B) (NS: M = 5.45, SD = 0.89; Baseline: M = 3.50, SD = 1.06, p = .002). $P_{7}$ found that ”turning exploration into a report feels almost effortless,” and $P_{17}$ highlighted how the system ”transforms explored branches into a structured story with one click.” These ratings align with the behavioral efficiency measures, indicating that participants both experienced and recognized the workflow improvements provided by NS, supporting DG1’s goal of fluid transitions from vague inquiries to evidence-grounded insights.
@@ 7.3.1. Participants Explored More Alternatives and Factors
With NS, participants considered significantly more alternatives during their analysis (M = 8.75, SD = 3.03, p ¡ 0.001) compared to Baseline participants (M = 3.25, SD = 1.58). They also explored a greater breadth of available factors, examining on average 7.0 out of 8 possible factors (SD = 0.9), compared to Baseline participants who explored 5.3 factors (SD = 1.4). As $P_{8}$ explained, ”With NS I felt like I could easily try different takes on the data, instead of sticking to one angle.”
@@
$P_{15}$ emphasized how NS encouraged breadth across factors: ”The system had nudged me to look at other dimensions when I reflected, like community well-being alongside cost.” Together, these findings suggest that NS not only increased the number of alternatives participants explored but also helped them consider broader coverage of available factors, creating more opportunities to surface contrasting perspectives.
@@ 7.3.2. Participants Integrated More Evidence into Decision-Making
Beyond exploring more alternatives, NS participants integrated more of their exploration into final decisions. They incorporated significantly more factors into their decision-making process (M = 3.8, SD = 0.7) compared to Baseline participants (M = 2.5, SD = 0.7). This suggests that NS participants not only explored more broadly but also utilized their exploration more meaningfully. Participants emphasized that the more exhaustive exploration in NS translated into having more actionable options to weigh. As $P_{11}$ explained, ”Because I had looked at more angles, I actually had choices, I could go in a few different directions instead of just one.” Similarly, $P_{6}$ elaborated, ”Having more alternatives in [NS] meant I didn’t have to stick with my first idea […] I could line up different options and actually remember how they compared, which made it easier to choose from.”
@@
7.4. [RQ4]: Reasoning Traceability
We assessed whether NS improved users’ ability to trace and justify their analytical reasoning process. Participants using NS reported significantly greater confidence in defending the reasoning behind their insights (M = 5.90, SD = 1.25) compared to Baseline (M = 3.15, SD = 1.60, p ¡ .001). Participants described how timelines and provenance features supported justification. $P_{2}$ explained, ”The timelines show the whole process which guarantees transparency and provides relevant evidence.” $P_{13}$ noted, ”Since the insights were linked to evidence and captured in the timeline, it is pretty straightforward to justify my reasoning.” Similarly, $P_{17}$ emphasized, ”Since every captured insight is automatically linked back to its supporting data and visualizations, I would always have clear evidence to justify my claims.” $P_{20}$ summarized this benefit as, ”I felt more confident that I could defend my choices because I had pulled in evidence from different places, not just one metric.”
When asked about presenting conclusions to others, participants also reported higher average confidence with NS (M = 5.95, SD = 1.00) than with Baseline (M = 5.15, SD = 1.60). However, this difference was not statistically significant (p = .112). Participants ($N=13$) rated themselves as confident presenting with NS, citing features such as organized evidence, visible branches, and insight timelines, which serve as valuable sources of support. For instance, $P_{1}$ stated, ”With all evidence organized, branches visible, and a polished story generated, I’d feel confident explaining my analysis to anyone.” Similarly, $P_{10}$ noted, ”Every insight is backed by reasoning, and the story mode is also helpful; therefore, it will aid in justifying my recommendations.”
Among these same participants, ($N=9$) also rated themselves as confident presenting with Baseline. In these cases, confidence stemmed from their own skills rather than the tool itself: $P_{1}$ remarked, ”I feel really confident because once I make a presentation, I will be able to address the weak points in my insights and conclusions.” Similarly, $P_{12}$ explained, ”I’d be comfortable with the recommendations themselves, but less confident explaining the full path I took to get there since some of that context lives outside the tool.”
@@
Experts explained that lower ratings often reflected a failure to capture implied intent. While generated dashboards were usually relevant to the surface wording of the prompt, they sometimes missed the deeper analytical angle that vague phrasing implied. For example, $E_{3}$ noted a prompt about ”some neighbourhoods are unexpectedly safe” suggested an outlier-focused analysis, but the system produced generic averages and correlations. Similarly, $E_{5}$ explained that when asked about ”travel growth differences between countries” the system produced growth rate visualizations that ”It showed similar percentages across countries, but missed that 10% growth could mean millions in one place and only thousands in another.” $E_{2}$ described such cases as ”technically correct but superficial”. This implies that while the majority of outputs were judged as good matches to the prompts, a smaller subset still fell short by failing to respond to the implied meaning behind vague inquiries.